hhanan:

What is another way to write x4/3

hhanan:

I got \[\sqrt[3]{x4}\] but want to make sure I am correct

XioGonz:

@jhonyy9 What do you think?

Joe2019:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @hhanan I got \[\sqrt[3]{x4}\] but want to make sure I am correct \(\color{#0cbb34}{\text{End of Quote}}\) i think you are right

hhanan:

thank you

jhonyy9:

\[\sqrt[3]{x^{4}}\] correct